The first of the three numbers exceeds twice the second number by 4. The third number is twice the first. If the sum of the three numbers is 54, find the numbers
Answer by ikleyn(52908)

.
The first of the three numbers exceeds twice the second number by 4.
The third number is twice the first. If the sum of the three numbers is 54, find the numbers.
~~~~~~~~~~~~~~~~

2nd number is x

1st number is 2x+4.

3rd number is 2(2x+4) = 4x+8.


The sum of the three numbers is 

    x + (2x+4) + (4x+8) = 54.


Simplify and find x

    7x + 12 = 54

    7x      = 54 - 122 = 42

     x                 = 42/7 = 6.


ANSWER.  2nd number is 6.  1st number is 2*6+4 = 16.  3rd number is 2*16 = 32.


CHECK.  The sum is 6 + 16 + 32 = 54.   ! Correct !

Solved.